OpenStudy (anonymous):

what fraction is equivalent to 0.081

OpenStudy (mathstudent55):

The 1 at the end of the number is in which decimal place?

OpenStudy (anonymous):

hundreths?

OpenStudy (anonymous):

\[0.081=\frac{8.1}{100}=\frac{81}{1000}\]

OpenStudy (mathstudent55):

The zero to the right of the decimal point is tenths. The 8 is hundredths. The 1 is thousandths. So the number is 81 thousandths. Just write that as a fraction: \( \dfrac{81}{1000} \)

OpenStudy (mathstudent55):

Since 81 = 3 * 3 * 3 * 3, and 1000 = 2 * 2 * 2 * 5 * 5 * 5, the two numbers have no common factors, so the fraction is reduced.

OpenStudy (anonymous):

You can break it down by resorting to going through the tens: \[0.081=\frac{0.081}{1}=\frac{0.81}{10}=\frac{8.1}{100}=\frac{81}{1000}\]

OpenStudy (anonymous):

alright sweet thanks sfor the help :)

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

Anytime!